OpenStudy (tootsi123):

Suppose that a cyclist began a 341 mi ride across a state at the western edge of the​ state, at the same time that a car traveling toward it leaves the eastern end of the state. If the bicycle and car met after 5.5 hr and the car traveled 37.2 mph faster than the​ bicycle, find the average rate of each.

OpenStudy (tootsi123):

The​ car's average rate is ___mph. ​(Type an integer or a​ decimal.) The​ bicycle's average rate is ____mph.

OpenStudy (thereisnousername):

Let cyclist speed be x mph car speed = (x+42.7)mph both of them take the same time to meet =5.5 hours distance: 341 miles combined speed: 2x+42.7 formula: t=d/r 5.5=341/(2x+42.7) 5.5(2x+42.7)=341 11x+85.4=341 11x=341-85.4 11x=255.6 x=23.24 Cycle speed, 23.24 + 42.7=65.94 Car speed 23.24= cycle speed 65.94= car speed

OpenStudy (tootsi123):

Thank you!

OpenStudy (tootsi123):

New question above. Please help.

OpenStudy (thereisnousername):

open a new question, so there isn't so much lag, and tag me

OpenStudy (tootsi123):

Okay I will try. :)